### v1

# Logical forms complement probability in understanding language model (and human) performance

Abstract
With the increasing interest in using large language models (LLMs) for planning in natural language, understanding their behaviors becomes an important research question. This work conducts a systematic investigation of LLMs’ ability to perform logical reasoning in natural language. We introduce a controlled dataset of hypothetical and disjunctive syllogisms in propositional and modal logic and use it as the testbed for understanding LLM performance. Our results lead to novel insights in predicting LLM behaviors: in addition to the probability of input (Gonen et al., 2023; McCoy et al., 2024), logical forms should be considered as orthogonal factors. In addition, we show similarities and differences between the logical reasoning performances of humans and LLMs by comparing LLM and human behavioral results.
Logical forms complement probability
in understanding language model (and human) performance
Yixuan Wang University of Chicago yixuanwang@uchicago.edu Freda Shi University of Waterloo Vector Institute, Canada CIFAR AI Chair fhs@uwaterloo.ca
1 Introduction
Logical reasoning is a fundamental aspect of building AI systems for reliable decision-making (Kautz et al., 1992, inter alia)—given a set of premises, an AI system should be able to deduce valid conclusions. With the advent of large language models (LLMs; Touvron et al., 2023; Jiang et al., 2023; AI@Meta, 2024, inter alia), there has been a surge of interest in using these models to assist planning and decision-making (Huang et al., 2022, inter alia); therefore, understanding the logical reasoning capabilities becomes crucial in understanding the reliability and potential of LLMs in planning. While recent work has shown that LLMs exhibit decent performance on logical reasoning problems (Liu et al., 2020; Ontanon et al., 2022; Wan et al., 2024, inter alia), there is still a lack of fine-grained understanding of the logical forms—among many argument forms presented in natural language (Shieber, 1993), do LLMs perform equally well, or do they exhibit preferences for certain argument forms? Do more complex components of logical forms, such as modalities, matter for LLM performance?
In this work, we investigate the logical reasoning capabilities of LLMs by assessing their performance on different logical forms. We curate a dataset of hypothetical syllogisms and disjunctive syllogisms in propositional and modal logic, where some logical forms are semantically equivalent to others. These forms are then interpreted with meaningful real-world statements that mirror the reasoning in daily communication, or phrases that consist of nonsensical words (Figure 2).
We conduct a series of controlled experiments to analyze the performance of a set of LLMs on these questions. Although our findings generally align with those by Gonen et al. (2023) and McCoy et al. (2024), who suggest that LLMs excel on examples with high probability, our results indicate that logical form, including but not limited to modalities and argument forms, is a crucial complementary factor in predicting the performance of LLMs (Figure 1). Additionally, with meaningful real-world interpretations, we find that:
1.
LLMs are still far from being perfect in atomic-level propositional and modal logic reasoning.
2.
LLMs prefer an affirmative answer under the modality of possibility, whereas they prefer a negative answer under the modality of necessity.
3.
In line with the recent results on categorical syllogisms (Eisape et al., 2024), we verify on hypothetical and disjunctive syllogisms that LLMs achieve better performance on certain logical forms that humans perform well. However, some logical forms receive favor from LLMs, while the phenomena lack support from human intuition or human behavioral data.
This paper is structured as follows. After reviewing related literature (§ 2), we describe the dataset synthesis process (§ 3). We report the LLM reasoning results on the synthesized dataset (§ 4) and ground the LLM reasoning performance to human reasoning performance (§ 5). We conclude by discussing the implications of our results and the limitations of this work (§ 6).
2 Related Work
Logical reasoning benchmarks. Prior LLM logical reasoning benchmarks (Liu et al., 2020; Han et al., 2022, inter alia) focus on complex, multi-hop reasoning problems with manually annotated problems, making cross-problem comparisons challenging. Recent work has introduced benchmarks with synthesized natural-language questions using predefined logical formulas and substitution rules (Saparov and He, 2022; Saparov et al., 2023; Parmar et al., 2024; Wan et al., 2024, inter alia). Compared to them, our work uniquely incorporates modal logic, which has been largely unexplored in existing benchmarks—while Holliday and Mandelkern (2024) present a case study, our approach offers two key advances: controlled knowledge bias in logic interpretations (§ 3.3) and a more rigorous statistical evaluation framework (§ 4.1).
Propositional and modal logic reasoning in language models. Recent work has explored training and finetuning language models specifically for logical reasoning (Clark et al., 2021; Hahn et al., 2021; Tafjord et al., 2022). Our work differs in two key aspects: we evaluate general-purpose language models through prompting, a cost-efficient setup that has been widely adopted in recent years, and we focus on propositional and alethic modal logic rather than temporal (Hahn et al., 2021) or epistemic (Sileo and Lernould, 2023) logic.111We acknowledge that any logic that involves non-truth-functional operators, including but not limited to first-order logic, temporal logic, and epistemic logic, can be viewed as a modal logic; however, we adopt the most restrictive sense of modal logic (Ballarin, 2023) and use it interchangeably with alethic modal logic. Additionally, unlike studies comparing LLM and human performance on categorical syllogisms (Eisape et al., 2024, inter alia),222 We refer readers to Zong and Lin (2024) for a more comprehensive review of categorical syllogisms. we focus on hypothetical and disjunctive syllogisms with considerations of modality.
Human logic reasoning.
Research on human reasoning capabilities has informed studies of LLM logical reasoning. Eisape et al. (2024) compared LLM syllogistic reasoning with human behavior results (Ragni et al., 2019) under the framework of the Mental Models Theory (Johnson-Laird, 1983). Lampinen et al. (2024) found similar content effects in human and LLM reasoning, supporting the need to control for common-sense knowledge in benchmarks (§ 3.2). Belem et al. (2024) studied human and LLM perception of uncertainty at a lexical level. Compared to them, we focus on the modal logic reasoning process and contribute new behavioral data.
3 Dataset
We curate a dataset of natural-language multi-choice questions to measure the logical inference performance of LLMs (Figure 2). First, we use both propositional and modal logical forms as templates (§ 3.1). Then, we assign meanings (e.g., real-world interpretations or phrases with nonsensical words) to each propositional variable in the templates, and translate templates into natural-language Yes/No questions (§ 3.2).
3.1 Background: Propositional and Modal Logic
Propositional logic studies the relation between propositions. In this framework, each proposition is typically represented by a variable, and multiple propositions combine with logical connectives (e.g., and ) to form compound propositions.
In propositional logic, a proposition can be evaluated as either true or false; however, this system can be overly simplistic when dealing with the complexity of real-world events. Consider the statement Alice is not eating., while it is true in a world where Alice is not eating, it may become false in a hypothetical possible world where Alice is indeed eating. This idea, known as possible world semantics (Kripke, 1959), provides a framework for expressing more nuanced statements about event possibilities, such as Alice may be eating and Alice must be eating. The former statement can be understood as there exists a possible world where Alice is eating, and the latter can be understood as in all possible worlds, Alice is eating.333 The possible world semantics, therefore, connects the notion of necessity and possibility to the universal and existential quantification () under first-order logic.
Normal modal logic (Kripke, 1963) follows this idea and extends propositional logic to reason about the necessity and possibility of events. A normal modal logic system can be formally described in the Backus–Naur form as
where is a propositional variable that serves as an atom in , is the negation operator, is the necessity operator (must), is the possibility operator (may), is logical disjunction (or), is logical conjunction (and), and is the logical implication operator (if…then). denotes the syntactic category of a formula in . The right-hand side of Eq. 1 describes all possible logical formulas under the system : for example, if , the rules imply that , , and so on. Following the convention in logic, the operator precedence is .
Indeed, the operators forms a functional complete set of operators under . Suppose and are variables that represent logical formulas. The logical or () and logical and () operators can be rewritten with logical not () and logical implication (), as follows:
Possibility operator can also be derived from the necessity operator.
Deduction and sequent.
Given a formula set as premises, if a deduction to a conclusion exists using axiom schemata and inference rules under the normal modal logic, we say the premises infer the conclusion, and the deduction can be represented as a logic sequent . If a formula set do not infer the conclusion, we denote it as and call it a non-entailment.
3.2 Translating Logic to Natural Language
An interpretation is a mapping from propositional variables to concrete meanings. For example, under the interpretation that is “Jane is eating apples” and is “John is eating oranges”, the logical formula becomes “Jane is eating apples or John is eating oranges.”
Choices of interpretation, i.e., the concrete content of the sentence, should not affect the underlying logical reasoning process. However, in natural-language utterances, reasoning can be influenced by various confounding factors. Knowledge bias is a common pitfall. For example, given the logical form , regardless of ’s interpretation, if we interpret then the conclusion will be “It is certain that cats are not animals.” But common-sense knowledge suggests that “it is certain that cats are animals” (), which logically contradicts the existing premise set.444 This confounding factor affects the examples in Table 18 of Han et al. (2022). Such bias will complicate logical reasoning (Lampinen et al., 2024) and should be avoided in data curation. Besides, each variable should have independent interpretation, as detailed in § B.1.
After being assigned interpretations, each logical form is further articulated as a yes-no question on whether the conclusion can be inferred from the premises. To mitigate the ambiguity in natural language, we design heuristic rules to translate logic forms into less ambiguous English, which are detailed in § B.1. For the exact wordings we used, see Table A1 in Appendices. If a valid deduction exists () for the logical form, the ground truth answer is Yes, otherwise No. The answer is solely determined by the logical form and is independent of the interpretation.
3.3 Involved Logical Forms
Translated logical forms can have varying degrees of naturalness. For example, the necessitation rule , which translates to “ is true; therefore, it is certain that is true,” appears to be unnatural due to redundancy. 555Nevertheless, we report the experiment results on necessitation rule in § C.1. Based on the relationship between and in Eq. 2, we use hypothetical and disjunctive syllogisms with four basic variants:
| (; modus ponens) | ||||
| (; modus tollens) |
Despite the semantic similarity, these logical forms translate to different natural-language questions. For example, taking the interpretations of and , translates to
| Consider the following statements: |
| Jane is watching a show or John is reading a book. |
| Jane isn’t watching a show. |
| Question: Based on these statements, can we infer that John is reading a book? |
With the same interpretation, translates to
| Consider the following statements: |
| If Jane isn’t watching a show, then John is reading a book. |
| Jane isn’t watching a show. |
| Question: Based on these statements, can we infer that John is reading a book? |
According to the commutativity of disjunction operator, we group and together as disjunctive syllogism, alongside two hypothetical syllogism groups, modus ponens () and modus tollens (). All the logical forms shown above are valid sequents with ground-truth answer Yes. To balance the dataset, we introduce some logic fallacies that generate questions with ground-truth label No. By flipping the second premises and the conclusions, we obtain the following fallacies:
where and are grouped as affirming the disjunction, and corresponds to affirming the consequent and denying the antecedent, respectively. In our dataset, we require the formulas and to the form of and , where and are propositional variables, each assigned with an interpretation. Both variables are constrained under the same modality , which can be necessity (), possibility () or no modality (). Pairing with four rules and theorem–fallacy variations, we have a total of forms.
3.4 Involved Logic Interpretations
For logic interpretations, we generate a set of verb phrases by prompting the CodeLlama 2 model (Rozière et al., 2024), and select 204 of them manually. and combine them with top-200 popular baby names in the US into subject-verb-object pairs,666https://www.ssa.gov/oact/babynames/names.zip such as . We randomly generate interpretations with two pairs each. The same set of interpretations is applied to variables in each logic sequent’s natural langauge template. In total, there are question, with samples shown in Table A1.
4 Experiment
4.1 Metrics and Investigated Models
Hu and Levy (2023) have suggested that the standard approach of greedily decoding yes-no strings (Dentella et al., 2023) may underestimate the competence of a language model; therefore, we adopt a probability-based metric to evaluate the model performance. In our evaluation protocol, the predicted likelihood of the tokens Yes and No, conditioned on the prompt —denoted as and , respectively—serve as the soft labels for yes-no answers. The soft accuracy on the single example with ground-truth answer is defined as the relative probability of :
where is the indicator function that returns 1 if the condition is true and 0 otherwise. This relative probability can also be viewed as the confidence score of the model on the ground-truth answer. The soft accuracy of a model on the entire dataset is defined as the average soft accuracy over all examples,
We use a zero-shot setting to investigate the general performance of the models’ logical inference capabilities—while adding detailed instructions or few-shot demonstrations may increase the absolute performance, they are at the cost of introducing possibly undesired confounding factors or behaviors, such as simply copy-pasting the answers in the examples.
We evaluate on the following models with open-sourced weights: mistral-7b-v0.2 and -8x7b (Jiang et al., 2023, 2024); llama-2-7b, -13b and -70b (Touvron et al., 2023); 3.1 version of llama-3-8b and -70b (AI@Meta, 2024); yi-34b (01.AI, 2024); phi-2 and phi-3-mini (Microsoft, 2023, 2024).777 Our evaluation protocol technically requires the conditional probabilities of specified answers given a prompt, which are not supported by most commercial models; however, we report the greedy-decoding accuracy of these models on a sample subset for reference.
4.2 Results: Performance w.r.t. Logical Forms
| Overall | Leaderboard | Modality | Argument Form | ||||||||||
| Model | (Rank) | (Rank) | |||||||||||
| mistral-7b | 0.645 | (4) | 0.145 | (7) | 0.464 | 0.496 | 0.974 | 0.877 | 0.663 | 0.280 | 0.434 | 0.653 | 0.939 |
| mistral-8x7b | 0.724 | (1) | 0.193 | (5) | 0.698 | 0.601 | 0.874 | 0.963 | 0.873 | 0.023 | 0.757 | 0.648 | 0.813 |
| llama-2-7b | 0.335 | (10) | 0.094 | (10) | 0.262 | 0.207 | 0.538 | 0.444 | 0.147 | 0.315 | 0.208 | 0.451 | 0.468 |
| llama-2-13b | 0.513 | (9) | 0.110 | (9) | 0.488 | 0.362 | 0.688 | 0.418 | 0.581 | 0.393 | 0.631 | 0.436 | 0.591 |
| llama-2-70b | 0.611 | (5) | 0.127 | (8) | 0.616 | 0.471 | 0.746 | 0.446 | 0.845 | 0.518 | 0.775 | 0.389 | 0.694 |
| llama-3-8b | 0.565 | (6) | 0.239 | (3) | 0.598 | 0.460 | 0.639 | 0.526 | 0.470 | 0.332 | 0.664 | 0.625 | 0.716 |
| llama-3-70b | 0.714 | (2) | 0.362 | (1) | 0.745 | 0.554 | 0.843 | 0.606 | 0.773 | 0.515 | 0.882 | 0.661 | 0.788 |
| yi-34b | 0.518 | (8) | 0.226 | (4) | 0.457 | 0.413 | 0.683 | 0.346 | 0.498 | 0.205 | 0.685 | 0.638 | 0.737 |
| phi-2 | 0.532 | (7) | 0.155 | (6) | 0.469 | 0.456 | 0.673 | 0.670 | 0.757 | 0.522 | 0.365 | 0.402 | 0.510 |
| phi-3-mini | 0.690 | (3) | 0.272 | (2) | 0.657 | 0.536 | 0.877 | 0.839 | 0.974 | 0.475 | 0.664 | 0.462 | 0.604 |
| OpenAI-o1 | 0.926 | N/A | N/A | N/A | 1.000 | 0.773 | 1.000 | 0.895 | 1.000 | 0.775 | 0.919 | 1.000 | 1.000 |
| Gemini-1.5-Pro | 0.859 | N/A | N/A | N/A | 0.831 | 0.748 | 0.997 | 1.000 | 1.000 | 0.919 | 0.661 | 0.991 | 0.638 |
| human | 0.595 | N/A | N/A | N/A | 0.589 | 0.566 | 0.640 | 0.691 | 0.901 | 0.628 | 0.594 | 0.225 | 0.411 |
We evaluate the aforementioned models with the probability-based protocol (Table 1). Generally, models that rank higher in the leaderboard also achieve higher soft accuracy on our dataset. The break-down accuracies on modalities and argument forms reveal that:
1.
(Modality) All models consistently perform better on the possibility () than necessity () or plain propositional logic.
2.
(Argument Forms) The pattern is more diverse, yet most of the models struggle the most on modus tollens () within logic sequents (i.e., questions with ground-truth answers Yes), and affirming the consequent () within fallacies.
| Hypothesis | -value |
4.2.1 Analysis on Logic Sequents
To systematically analyze the effect on model performance of each factor of interest, as well as cross-validating the observations above, we fit a linear mixed-effects model (Raudenbush, 2002) to the soft accuracy data on valid logic sequents (i.e. with ground truth of Yes) across different LLMs and logical forms,
with the linear fixed effects of (i.) modality, (ii.) argument form, and (iii.) input perplexity. Individual probability, coupled with a constant term, is modeled as a random effect to account for potential model-specific biases. Here, Perplexity denotes the perplexity of the input text (), which is defined as the exponential of the token-wise average negative log-likelihood of the text given a specific language model:
The mixed-effects model yields a marginal of and a conditional of , suggesting a reasonable predictive power. The likelihood ratio test on the full regression model vs. the null regression model without each of the fixed effects yields a significant result (), suggesting the importance of all these factors in determining the model performance.
Fixed effects. In line with Gonen et al. (2023) and McCoy et al. (2024), we find a negative correlation between perplexity and soft accuracy (); however, the correlation between them is weak (), which suggests the necessity of the complementary factors below in predicting LLM performance.
For different modalities and argument forms, we estimate their marginal means on soft accuracy (Figure 3), and perform pairwise hypothesis testing on the estimated coefficients (Table 2). The results generally align with the general observations on the full dataset. The only exception is that modus ponens (), instead of disjunctive syllogism (), appears to be the easiest argument form (i.e., the one with the highest soft accuracy) among all.
Random effects. We analyze the per-LLM random effects on the soft accuracy (Figure 4). All the model-specific mixed effects of perplexity are negative, suggesting the negative correlation between perplexity and soft accuracy is consistent across models (Figure 4a). While the intercept random effects are not perfectly aligned with the model performance—since the perplexity random effects may introduce confounding factors—higher-ranked models generally tend to have higher intercept random effects (Figure 4b), which cross-validates the general performance ranking.
4.2.2 Extended Analysis on the Negative Perplexity–Performance Correlation
We further investigate the negative correlation between perplexity and model performance through a controlled experiment: we create a mirror dataset of the same size, keeping all the logical formulas while interpreting them with nonsensical words. For example, the formula may be interpreted as it’s possible that Neva is balaring a montery or Lucille is sweeling prandates, where the underlined words and phrases are nonsensical. Intuitively, the perplexity of the problems in this mirror dataset should be much higher than that of the primary dataset problems (§ 3) under any reasonably trained language model.
We analyze the correlation between perplexity and model performance (Figure 5). As desired, the perplexity of problems with nonsensical words are indeed much higher than that of the primary dataset ( vs. ). The significant portion of horizontal and inclined lines in the figures again suggests that perplexity is not a reliable predictor of model performance. Meanwhile, the overall parallelism of the lines echos our results that logical forms are important factors for such prediction.
4.2.3 The Affirmation Bias over Modalities
One key argument of Dentella et al. (2023) is that large language models exhibit a bias towards affirming the claim, i.e., answering Yes more frequently than No.
We investigate this phenomenon by fitting a mixed-effects model
which has the same structure as Eq. 4, except the dependent variable being the relative probability of answering Yes conditioned on input text .
We present the estimated marginal means of the factors in the mixed-effects model (Figure 6). While our results confirm the affirmation bias on propositional logic, such bias is slightly less pronounced on the possibility modality (, around 0.03), and the models even show a bias towards rejecting claims under the necessity modality ().
5 Human Experiments
LLMs are trained on text produced by humans and are able to generate plausible text; therefore, there have been interests in using LLMs as human models (Eisape et al., 2024; Misra and Kim, 2024, inter alia). Following this line of work, we conduct a human behavioral experiment to ground the LLM reasoning behavior. Using samples from our primary dataset, we collected 710 responses from adults fluent in English through Prolific.888https://prolific.com More experiment details can be found in § A.2.
The average human accuracy on each group is shown in the last row of Table 1.999Human responses are binary classes, so correct and incorrect responses are coded as and , respectively. Aligned with our LLM results (§ 4), on modalities the overall human results also show an accuracy order of (), and on argument forms, modus ponens () is the most accurately answered pattern.
To further investigate the interactions of logic factors, we fit a generalized linear mixed-effects model (Bates et al., 2015) to verify the effect of modality and argument forms on human logic reasoning accuracy (Eq. 6 and Figure 7).
where is the binary accuracy of human responses, and Rt is the response time. The generalized mixed-effects model yields a marginal of yet a conditional , indicating a diverse response pattern across participants. The likelihood ratio test on the full model against the null model shows that only the effect of argument form is significant (, ). However, in accordance with the overall performance, we find modus ponens () has a significantly higher effect than other two valid argument forms. This confirms that logical forms can also have a significant impact on human reasoning accuracy, which is consistent with the LLM results, although the effect sizes are not the same.
6 Conclusion and Discussion
We present an analysis of hypothetical and disjunctive syllogisms on propositional and modal logic and systematically analyze the LLM performance on the dataset. Our analysis provides novel insights on explaining and predicting LLM performance: in addition to the perplexity or probability of the input text, the underlying logic forms play an important role in determining the performance of LLMs. In addition, we compare the behaviors of LLMs and humans using the same data through human behavioral experiments. We discuss the implications of our results as follows.
Probability in language models.
Probability and, relatedly, perplexity are often used as intrinsic evaluation metrics for language models. While Gonen et al. (2023) and McCoy et al. (2024) show that probability and perplexity, to some extent, correlates with LLM performance, literature in program synthesis with LLMs shows little correlation between probability and execution-based evaluation results (Li et al., 2022; Shi et al., 2022). This work does not necessarily contradict either line, but rather provides complementary factors for analyzing LLM performance.
We argue that probability may have become an overloaded term in analyzing LLMs. Low probability may be due to one or more of the following non-exhaustive reasons: (1) the sentence is out of context, (2) the sentence is in ungrammatical language, or (3) the sentence is grammatical but has semantically awkward content (e.g., the mirror dataset in § 4.2.2), (4) the sentence contains rare content while being reasonable. We hypothesize that the probability of language models may not be essentially able to capture all these nuanced differences, and call for encoding and decoding algorithms—such as Meister et al. (2023)—that can better decompose the probability into finer-grained components.
Comparing humans and LLMs.
What is our goal for building LLMs? Is it to achieve better performance on practical tasks or to build a more human-like model? Our results, together with Eisape et al. (2024), suggest that these two goals may not be perfectly aligned by revealing a mixture of similarity and discrepancy between LLMs and humans—for example, while LLMs exhibit higher benchmark performance than humans on our dataset and show the same argument form preferences with humans (Figures 3 and 7), they also show systematic biases that humans do not have (e.g., disfavoring the necessity modality, § 4.2.3). While there has been positive evidence of using LLMs as human models in psycholinguistic studies (Misra and Kim, 2024, inter alia), our results reaffirm that we should execute such approaches cautiously.
On the relation between modality and performance.
Our results show that there is a significant difference in performance between necessity and possibility modalities, with the former much lower than the latter (Table 1). Part of the reason for this is that LLMs have a significant tendency to say “No” to necessity modality (Figure 6).
On the one hand, our results extend the conclusion of Dentella et al. (2023), who found that LLMs generally respond positively—LLM behaviors may be significantly affected by finer-grained factors, including but not necessarily limited to the logical modality of the input. On the other hand, while LLMs systematically tend to answer “No” to questions in necessity modality, we do not find related evidence in human experiments, which leads to us hypothesize that such rejection bias comes from either the model architecture or the training strategies, such as the reinforcement learning with human feedback (RLHF; Ouyang et al., 2022) protocol. We leave this as an open question for future research.
Modal logic and theory of mind.
Modality, in principle, encodes mental states and beliefs. The reasoning of beliefs also resonates with the theory of mind (Premack and Woodruff, 1978; Baron-Cohen et al., 1985, inter alia) and machine theory of mind (Rabinowitz et al., 2018; Ma et al., 2023, inter alia). Following the effort by Sileo and Lernould (2023) that uses epistemic modal logic to model the machine theory of mind, our work assesses the behaviors of LLMs on alethic modal logic, distantly revealing the future potential of LLMs in achieving the theory of mind.
Limitations
This work comes with two major limitations:
1.
While we have verified that our data has a low perplexity ( under mistra-7b; much lower than that of the data by Wan et al. (2024), ), and, therefore, are similar enough to natural language utterances, the synthetic language cannot fully substitute natural language in daily life. Our dataset and analysis are not comprehensive enough to cover many nuanced examples that may appear in real communication, especially when context-dependent understanding is crucial to conveying communication goals.
2.
Despite more than 7,000 languages worldwide, as a first step, our material only covers English. This narrow focus is due to the languages the authors are proficient in and the coverage of the language models. We acknowledge the importance of extending the scope of this work to a more comprehensive set of languages and leave the extension as an immediate follow-up step.
In addition, the sample size of human experiments is somewhat limited. We leave more comprehensive human behavioral data collection and analysis to future work.
Ethics Statement
While this work involves human logical reasoning experiments, we have ensured that (1) the data are generated procedurally following templates listed in the paper and (2) there is no harmful content in the atomic logical interpretations, reviewed by all the authors. In addition, we have ensured that all participants are paid a fair wage through the Prolific platform. Instructions and consent forms delivered to the participants can be found in the § A.2. The institutional ethics review board has approved the data collection process.
We do not foresee risk beyond the minimal risk posed by LLM evaluation work. We acknowledge that using LLMs in real-world scenarios could significantly impact human behaviors, raising the need for model transparency, safety, security, and interpretability. This work contributes to the understanding of We will open-source the synthetic logical reasoning dataset upon publication.
7 Acknowledgements
We thank Yudong Li for his help in setting up the Gemini and OpenAI API for the experiments. This work was supported in part by a Google PhD Fellowship and a Canada CIFAR AI Chair award to FS, as well as NSERC RGPIN-2024-04395.
Appendix A Additional Experiment Details
A.1 LLM Experiment Details
All LLMs used are obtained from Hugging Face checkpoints. Time and compute power requirements vary, the largest llama-3-70b model takes around 2 hours on NVIDIA A6000 GPU to obtain all results in § 4.
A.2 Human Experiment Details
Participant consent.
We use the following language to obtain consent from participants, where our institution name is replaced with the Anonymous Institution to protect the anonymity of submission.
This study is part of a scientific research project at the Anonymous Institution. Your decision to complete this study is voluntary. There is no way for us to identify you. The only information we will have, in addition to your responses, is the demographic information you provided to Prolific and the time at which you completed the survey. The results of the research may be presented at scientific meetings or published in scientific journals. Clicking on the button below indicates that you are at least 18 years of age and agree to complete this study voluntarily. Press the button below to start the experiment.
Participant instructions.
We use keys F and J, which are roughly symmetric on a standard English keyboard, to collect participant responses. Half of the participants see the following instruction:
In this study, you will be presented with two statements followed by a question. Your task is to answer either Yes or No to the question, based on the information provided in the statements. Please respond quickly and accurately by pressing "F" for Yes, and "J" for No.
To mitigate the possible bias introduced by the dominant hand, we have the other half of the participants see instruction with reversed keys: In this study, you will be presented with two statements followed by a question. Your task is to answer either Yes or No to the question, based on the information provided in the statements. Please respond quickly and accurately by pressing "F" for No, and "J" for Yes.
Participant wage.
We offer participants an hourly wage of 1.5 times Prolific’s minimum wage. The duration is determined by the median completion time among all participants.
Appendix B Extra Details of the Dataset
B.1 Considerations in Translating Logical Form to Natural Language
During the interpretation process, another key point is to assign independent interpretations to variables. Deciding the dependency also involves common sense knowledge. For example, consider the premises and . If we interpret and to proposition variables and , the two variables are possibly not independent. According to common sense, “Jane is not inside the house” () correlates with or is even equivalent to “Jane is out” (). Logically, ; however, with the extra premise given by common sense, people may conclude that .101010 This confounding factor affects the examples in Figure 10 of Holliday and Mandelkern (2024).
Besides, natural language is ambiguous—one sentence in natural language can come from multiple logical forms under the same interpretation. We use present tense and progressive aspect to encourage a reading of imaginary ongoing events, corresponding to the alethic modality. Such events are less likely to induce LLM’s or human’s individual bias, as they are unrelated to factual knowledge or moral judgements. Also, we always use two full verb phrases, ruling out sentences like “Jane is eating apples or oranges,” so the two events are less likely to be mutually exclusive. In this way, we can reduce the ambiguity of the questions in our dataset.
B.2 Data Samples
All logic forms and corresponding natural language sentences can be found in Table A1.
The exact prompt format is as follows:
[…]
Answer:<eof>
| Validity | Modality | Argument Form | Logical Form | Natural Language |
Jane is watching a show or John is reading a book.
Jane isn’t watching a show. Can we infer that John is reading a book? |
Jane is watching a show or John is reading a book.
John isn’t reading a book. Can we infer that Jane is watching a show? |
If Jane isn’t watching a show, then John is reading a book.
Jane isn’t watching a show. Can we infer that John is reading a book? |
If Jane isn’t watching a show, then John is reading a book.
John isn’t reading a book. Can we infer that Jane is watching a show? |
It’s certain that Jane is watching a show or it’s certain that John is reading a book.
It’s uncertain whether Jane is watching a show. Can we infer that it’s certain that John is reading a book? |
It’s certain that Jane is watching a show or it’s certain that John is reading a book.
It’s uncertain whether John is reading a book. Can we infer that it’s certain that Jane is watching a show? |
If it’s uncertain whether Jane is watching a show, then it’s certain that John is reading a book.
It’s uncertain whether Jane is watching a show. Can we infer that it’s certain that John is reading a book? |
If it’s uncertain whether Jane is watching a show, then it’s certain that John is reading a book.
It’s uncertain whether John is reading a book. Can we infer that it’s certain that Jane is watching a show? |
It’s possible that Jane is watching a show or it’s possible that John is reading a book.
It’s impossible that Jane is watching a show. Can we infer that it’s possible that John is reading a book? |
It’s possible that Jane is watching a show or it’s possible that John is reading a book.
It’s impossible that John is reading a book. Can we infer that it’s possible that Jane is watching a show? |
If it’s impossible that Jane is watching a show, then it’s possible that John is reading a book.
It’s impossible that Jane is watching a show. Can we infer that it’s possible that John is reading a book? |
If it’s impossible that Jane is watching a show, then it’s possible that John is reading a book.
It’s impossible that John is reading a book. Can we infer that it’s possible that Jane is watching a show? |
Jane is watching a show or John is reading a book.
John is reading a book. Can we infer that Jane isn’t watching a show? |
Jane is watching a show or John is reading a book.
Jane is watching a show. Can we infer that John isn’t reading a book? |
If Jane isn’t watching a show, then John is reading a book.
John is reading a book. Can we infer that Jane isn’t watching a show? |
If Jane isn’t watching a show, then John is reading a book.
Jane is watching a show. Can we infer that John isn’t reading a book? |
It’s certain that Jane is watching a show or it’s certain that John is reading a book.
It’s certain that John is reading a book. Can we infer that it’s uncertain whether Jane is watching a show? |
It’s certain that Jane is watching a show or it’s certain that John is reading a book.
It’s certain that Jane is watching a show. Can we infer that it’s uncertain whether John is reading a book? |
If it’s uncertain whether Jane is watching a show, then it’s certain that John is reading a book.
It’s certain that John is reading a book. Can we infer that it’s uncertain whether Jane is watching a show? |
If it’s uncertain whether Jane is watching a show, then it’s certain that John is reading a book.
It’s certain that Jane is watching a show. Can we infer that it’s uncertain whether John is reading a book? |
It’s possible that Jane is watching a show or it’s possible that John is reading a book.
It’s possible that John is reading a book. Can we infer that it’s impossible that Jane is watching a show? |
It’s possible that Jane is watching a show or it’s possible that John is reading a book.
It’s possible that Jane is watching a show. Can we infer that it’s impossible that John is reading a book? |
If it’s impossible that Jane is watching a show, then it’s possible that John is reading a book.
It’s possible that John is reading a book. Can we infer that it’s impossible that Jane is watching a show? |
If it’s impossible that Jane is watching a show, then it’s possible that John is reading a book.
It’s possible that Jane is watching a show. Can we infer that it’s impossible that John is reading a book? |
Appendix C Additional Experiments
C.1 Extra Experiment: Introduction Rule of Modality
We report the results on the necessitation rule and its variants here, as these rules are obscure and verbose to be articulated in natural language:
| (necessitation rule) | ||||
Its natural language form is as follows:
| Jane is watching a show. | |
| () | Can we infer that it’s certain that Jane is watching a show? |
| () | Can we infer that it’s possible that Jane is watching a show? |
| () | Can we infer that Jane is watching a show? |
All three variants are paired with 1000 logic interpretations. As they are all rules of inference, the ground truth answer is always Yes. Overall accuracy is shown in Table A2, where across all LLMs, the necessitation rule has the lowest accuracy. This echoes the necessity modality’s tendency to be rejected discussed in § 4.2.3.
| mistral-7b | 0.998 | 0.885 | 0.999 |
| mistral-8x7b | 0.957 | 0.540 | 0.987 |
| llama-2-7b | 0.768 | 0.013 | 0.920 |
| llama-2-13b | 0.368 | 0.004 | 0.829 |
| llama-2-70b | 0.511 | 0.051 | 0.834 |
| llama-3-8b | 0.398 | 0.225 | 0.783 |
| llama-3-70b | 0.674 | 0.384 | 0.794 |
| yi-34b | 0.960 | 0.382 | 0.999 |
| phi-2 | 0.814 | 0.226 | 0.892 |
| phi-3-mini | 0.992 | 0.925 | 0.994 |
We further fit a linear mixed-effects model similar to Eq. 4, except that the argument form effect is now constant across all data points. The mixed-effects model yields a marginal of and a conditional of . Estimated marginal means shows that the accuracy on is less than , but higher than , with both differences significant at . This further suggests that modality serves as an important factor on logic reasoning performance.
C.2 Extra Experiment: Distribution of Modalities
Besides the necessitation rule, distribution axiom is the other fundamental axiom in normal modal logic. It can be transformed into the rule shown in Eq. A1, and plugging in the definition of in Eq. 2 gives the rule shown in Eq. A2. Notice that Eq. A2 closely resembles rule ’s variant with necessity, as shown in Eq. A3, except the different scope of the necessity operator and the position of the negation operator. Moving the negation operator out of the necessity operator will result in a fallacy (Eq. A4).
We say A2, A3 and A4 are of argument form theorem, base and spurious, respectively. See Table A3 for the logical forms and their ground truth we used to study the distribution of modalities. The natural language form is as follows:
| It’s certain that if Freddy is not going shopping, then Coy is making dinner. | |
| (theorem) | It’s certain that Freddy is not going shopping. |
| (spurious) | It’s uncertain whether Freddy is going shopping. |
| Can we infer that it’s certain that Coy is making dinner? |
This group of rules and fallacies comes from the fact that the necessity modality is not distributive to disjunction, i.e. (Xiang, 2019, Ex. 5). In contrast, the possibility modality is distributive to disjunction. This particular case could have served as a material to test the LLM’s knowledge of the asymmetry between the two modalities, yet in § 4.2.3 we showed that there is a bias towards rejection on the necessity modality. As the false case of the disjunction is on the necessity modality, this bias confounds the experiment.
| Modality | Argument Form | Logical Form |
| base | ||
| base | ||
| theorem | ||
| spurious | ||
| base | ||
| theorem | ||
| spurious |
We fit a linear mixed-effects model similar to Eq. 4 to the data,
with an interaction term between the modality and argument form. On the theorem form compared to the base form, the necessity modality has a higher estimated marginal means with significance, yet the possibility modality has a lower estimated marginal means. On the spurious form compared to the base form, the has a higher means, and the has no significant difference. On both forms, in terms of accuracy still holds at a slight margin of and respectively.
To verify whether on the performance increase on spurious form is due to the rejection bias, we fit a linear mixed-effects model with the relative probability of answering Yes as dependent variable. Results show that on spurious form compared to the base form, the effect of ’s tendency to answer Yes is only lower, indicating the rejection bias of the base form is still present. Therefore, we hypothesize that the LLM’s performance on recognizing the fallacy of necessity distribution over disjunction is hindered by the rejection bias on the necessity modality.